OpenStudy (anonymous):

14x+18y=-54 -4x-14y=42 Solve using the elimination method

OpenStudy (anonymous):

\[-56x-72y=216\]\[-56x-196y=588\] Subtract the lower from the upper: \[124y=-372\]\[y=-3\] plug y into the first:\[14x-54=-54\]\[x=0\]